AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 298     Entry time: Sun Feb 27 06:58:51 2022
Author: RSS TD 
Subject: Current status 

7:59: 95K-->120K

8:06: 120K --> 95K

Ring current ~7e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.44 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 8-xy-2022-02-27_08-10-15.png  275 kB  | Hide | Hide all
8-xy-2022-02-27_08-10-15.png
Attachment 2: 8-ey-2022-02-27_08-09-46.png  168 kB  | Hide | Hide all
8-ey-2022-02-27_08-09-46.png
Attachment 3: 8-ex-2022-02-27_08-09-29.png  166 kB  | Hide | Hide all
8-ex-2022-02-27_08-09-29.png
Attachment 4: Screenshot_from_2022-02-27_07-16-30.png  271 kB  | Hide | Hide all
Screenshot_from_2022-02-27_07-16-30.png
ELOG V3.1.4-unknown